# and Aggregation Improves Adversarial Robustness

### v5

@@ 6 Conclusion
In this work, we presented a novel adversarial defence mechanism tightly integrating input purification, and classification by an adversarially-trained model – in the form of representation-conditional data purification. Our method is able to improve upon the current state-of-the-art in Cifar-10, Cifar-100, and TinyImageNet $\ell_{\infty}$ robust classification, w.r.t. both adversarial training and purification approaches alone.
@@ Acknowledgements
Appendix A On Projected Gradient Descent adversarial training
The task of determining model parameters $\bm{\theta}^{\star}$ that are robust to adversarial perturbations is cast in (Madry et al., 2018) as a min-max optimisation problem seeking to minimise adversarial risk, i.e.:
$\bm{\theta}^{\star}\approx\hat{\bm{\theta}}^{\star}:=\operatorname*{\arg\min}_{\bm{\theta}}\mathbb{E}_{\left(\bm{x},y\right)\sim\mathcal{D}}\left[{\max_{\bm{\delta}\in\mathbb{S}}\mathcal{L}\left({f\left({\bm{x}+\bm{\delta};\bm{\theta}}\right),y}\right)}\right]$
where $\mathcal{D}$ is the distribution on the examples $\bm{x}$ and the corresponding labels $y$, $f(\cdot;\bm{\theta})$ is a model with learnable parameters $\bm{\theta}$, $\mathcal{L}$ is a suitable loss function, and $\mathbb{S}$ is the set of allowed constrained perturbations. In the case of $\ell_{p}$ norm-bound perturbations of maximum magnitude $\epsilon$, we can further specify $\mathbb{S}:=\left\{{\bm{\delta}\operatorname*{|}\|\bm{\delta}\|_{p}\leq\epsilon}\right\}$.
The inner optimisation problem is solved, by Madry et al. (2018), by Projected Gradient Descent (Pgd), an iterative algorithm whose goal is the synthesis of an adversarial perturbation $\hat{\bm{\delta}}=\bm{\delta}^{\left(K\right)}$ after $K$ gradient ascent and projection steps defined as:
$\bm{\delta}^{\left(k+1\right)}\leftarrow\mathfrak{P}_{\mathbb{S}}{\left({\bm{\delta}^{\left(k\right)}+\alpha\operatorname{sign}{\left({\nabla_{\bm{\delta}^{\left(k\right)}}{\mathcal{L}_{ce}({{f({\bm{x}+\bm{\delta}^{\left(k\right)};\bm{\theta}}),y})}}}\right)}}\right)}$
where $\bm{\delta}^{\left(0\right)}$ is randomly sampled within $\mathbb{S}$, $\alpha$ is a hyperparameter (step size), $\mathcal{L}_{ce}$ is the cross-entropy function, and $\mathfrak{P}_{\mathbb{A}}$ is the Euclidean projection operator onto set $\mathbb{A}$, i.e.:
$\mathfrak{P}_{\mathbb{A}}(\bm{a}):=\operatorname*{\arg\min}_{\bm{a}^{\prime}\in\mathbb{A}}\left|\left|\bm{a}-\bm{a}^{\prime}\right|\right|_{2}\text{ .}$
The outer optimisation is carried out by simply training $f(\cdot;\bm{\theta})$ on the examples found by Pgd against the current model parameters – and their original pre-perturbation labels. The overall procedure just described constitutes Pgd adversarial training.
Appendix B On the functioning of (conditional) Variational Autoencoders
Variational autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014) learn from data a generative distribution of the form $p(\bm{x},\bm{z})=p(\bm{x}\operatorname*{|}\bm{z})p(\bm{z})$, where the probability density $p(\bm{z})$ represents a prior over latent variable $\bm{z}$, and $p(\bm{x}\operatorname*{|}\bm{z})$ is the likelihood function, which can be used to sample data of interest $\bm{x}$, given $\bm{z}$.
Training is carried out by maximising a variational lower bound, $-\mathcal{L}_{\text{VAE}}(\bm{x})$, on the log-likelihood $\log p(\bm{x})$ – which is a proxy for the Evidence Lower Bound (ELBO) – i.e.:
$-\mathcal{L}_{\text{VAE}}(\bm{x}):=\mathbb{E}_{q(\bm{z}\operatorname*{|}\bm{x})}[\log p(\bm{x}\operatorname*{|}\bm{z})]-\operatorname{KL}(q(\bm{z}\operatorname*{|}\bm{x})\|p(\bm{z}))$
where $q(\bm{z}\operatorname*{|}\bm{x})\approx p(\bm{z}\operatorname*{|}\bm{x})$ is an approximate posterior and $\operatorname{KL}(\cdot\|\cdot)$ is the Kullback-Leibler divergence.
By parameterising the likelihood with a decoder ANN $p_{\bm{\theta}_{\text{D}}}(\bm{x}\operatorname*{|}\bm{z};\bm{\theta}_{\text{D}})\approx p(\bm{x}\operatorname*{|}\bm{z})$, and a possible variational posterior with an encoder ANN $q_{\bm{\theta}_{\text{E}}}(\bm{z}\operatorname*{|}\bm{x};\bm{\theta}_{\text{E}})\approx q(\bm{z}\operatorname*{|}\bm{x})$, the parameters $\bm{\theta}_{\text{D}}^{\star}$ of the generative model that best reproduces the data can be learnt – jointly with $\bm{\theta}_{\text{E}}^{\star}$ – as:
$\bm{\theta}_{\text{E}}^{\star},\bm{\theta}_{\text{D}}^{\star}:=$
$\operatorname*{\arg\min}_{(\bm{\theta}_{\text{E}},\bm{\theta}_{\text{D}})}\mathcal{L}_{\text{VAE}}(\bm{x})=$
$\operatorname*{\arg\min}_{(\bm{\theta}_{\text{E}},\bm{\theta}_{\text{D}})}\mathbb{E}_{\bm{x}\sim\mathcal{D}}\left[{-\mathbb{E}_{\bm{z}\sim{q_{\bm{\theta}_{\text{E}}}(\bm{z}\operatorname*{|}\bm{x};\bm{\theta}_{\text{E}})}}\left[{\log{p_{\bm{\theta}_{\text{D}}}(\bm{x}\operatorname*{|}\bm{z};\bm{\theta}_{\text{D}}})}\right]}+\operatorname{KL}({q_{\bm{\theta}_{\text{E}}}(\bm{z}\operatorname*{|}\bm{x};\bm{\theta}_{\text{E}})}\|{p(\bm{z})})\right]$
where $\mathcal{D}$ is the distribution over the (training) examples $\bm{x}$.
From a practical point of view, optimisation is based on the empirical evaluation of $\mathcal{L}_{\text{VAE}}(\bm{x};\bm{\theta})$ on mini-batches of data, with the term $-\mathbb{E}_{\bm{z}\sim{q_{\bm{\theta}_{\text{E}}}(\bm{z}\operatorname*{|}\bm{x};\bm{\theta}_{\text{E}})}}\left[{\log{p_{\bm{\theta}_{\text{D}}}(\bm{x}\operatorname*{|}\bm{z};\bm{\theta}_{\text{D}}})}\right]$ replaced by a reconstruction cost
$\mathcal{L}_{\text{Reco}}(\bm{x},\bm{x^{\prime}})\geq 0\operatorname*{|}\mathcal{L}_{\text{Reco}}(\bm{x},\bm{x}^{\prime})=0\iff x=x^{\prime}\text{ .}$
The generation of new data according to the fitted model is achieved by sampling from
$p_{\bm{\theta}_{\text{D}}^{\star}}(\bm{x}\operatorname*{|}\bm{z};\bm{\theta}_{\text{D}}^{\star})\biggr{\rvert}_{\bm{z}\sim p(\bm{z})}$
i.e. decoding samples from $p(\bm{z})$.
The setting is analogous in the case of conditional Variational Autoencoders (Sohn et al., 2015; Yan et al., 2016) (see section 3), where conditional sampling is achieved by
$\bm{x}_{\bm{c_{j}}}\sim p_{\bm{\theta}_{\text{D}}^{\star}}(\bm{x}\operatorname*{|}\bm{z},\bm{c};\bm{\theta}_{\text{D}}^{\star})\biggr{\rvert}_{\bm{z}\sim p(\bm{z});\ \bm{c}=\bm{c_{j}}}\text{ .}$
Appendix C Justification of Adversarially-balanced batches
@@
Appendix D Justification of the robust aggregation strategy
@@ D.1 Heuristic analysis
In the former scenario, the product of $e^{{l_{i}^{\alpha}}}$ terms can be arbitrarily deflated (up to zero) by lowering the $l_{i^{\star}}^{\alpha^{\star}}$ logit only. In the latter scenario, the logit can be arbitrarily inflated, and such effect is only partially suppressed by normalisation by $Z$ (a sum of $\nicefrac{{1}}{{N}}$-exponentiated terms).
@@
In the former scenario, the effect of logit deflation on a single product term is bounded by $e^{e^{l_{i^{\star}}^{\alpha^{\star}}}}>1$, thus exerting only a minimal collateral effect on the product, through a decrease of $Z$. This effectively prevents aggregation takeover by logit deflation. Similarly to logit averaging, in the latter scenario, the logit can be arbitrarily inflated. However, in this case, the effect of normalisation by $Z$ is much stronger, given its increased magnitude.
@@
D.2 Experimental analysis
To further corroborate the choice of the specific aggregation function described in subsection 4.5, we repeat the assessment of Carso under the same conditions described in section 5, the only difference being the use of the alternative aggregation functions analysed in subsection D.1 (i.e. logit and probability aggregation). Results, in terms of both clean and robust accuracy, are shown in Table 4.
As we can see, the use of the alternative aggregation strategies leads to a minimal variations in the clean accuracy attained, whereas the corresponding robust accuracy sharply decreases – in the case of Cifar-10 even below random chance – as the attacks become increasingly effective. Such results strongly corroborate the use of the normalised doubly-exponential logit product proposed as the aggregation strategy in subsection 4.5 and prove its central role in the overall adversarial robustness of the method.
@@ D.2 Experimental analysis
Appendix E Architectural details and hyperparameters
In the following section, we provide more precise details about the architectures (subsection E.1) and hyperparameters (subsection E.2) used in the experimental phase of our work.
E.1 Architectures
@@ E.2 Hyperparameters
Fgsm Pgd Pgd+EoT Input clipping $[0.0,1.0]$ $[0.0,1.0]$ $[0.0,1.0]$ of steps $1$ $40$ $200$ Step size $\epsilon_{\infty}$ $0.01$ $0.007$ Loss function CCE CCE CCE # of EoT iterations $1$ $1$ $20$ Optimiser SGD SGD
@@
All scenarios Sc. (a) Sc. (b) Sc. (c) Optimiser RAdam$+$Lookahead RAdam $\beta_{1}$ $0.9$ RAdam $\beta_{2}$ $0.999$ RAdam $\epsilon$ $10^{-8}$ RAdam Weight Decay None Lookahead averaging decay $0.8$ Lookahead steps $6$ Initial LR $5\times 10^{-9}$ Loss function CCE Sampled reconstructions per input $8$ Epochs $200$ $200$ $250$ LR warm-up epochs $25$ $25$ $31$ LR plateau epochs $25$ $25$ $31$ LR annealing epochs $150$ $150$ $188$ Plateau LR $0.064$ $0.064$ $0.0128$ Final LR $4.346\times 10^{-4}$ $4.346\times 10^{-4}$ $1.378\times 10^{-4}$ $\beta$ increase initial epoch $25$ $25$ $32$ $\beta$ increase final epoch $34$ $34$ $43$ Batch size $5120$ $2560$ $1024$ Adversarial batch fraction $0.5$ $0.15$ $0.01$
Appendix F Additional tables
@@
The following section contains additional tabular data that may be of interest to the reader.
Table 16 reports the respective clean accuracies for the best models available in terms of AutoAttack robust accuracy, in scenarios (a) and (b). Models are further divided in AT-based and purification-based, so as to match the corresponding columns for robust accuracy shown in Table 2.
The best AT-based model for Cifar-10 is taken from Cui et al. (2023), whereas that for Cifar-100 from Wang et al. (2023). Both best purification-based models are taken from (Lin et al., 2024).
The clean and robust accuracies for the best AT-based model on TinyImageNet-200 (scenario (c)) are already part of Table 2 and we redirect the reader there for such information. We are not aware of any published state-of-the-art adversarial purification-based model for TinyImageNet-200.
Table 16: Clean accuracy for the best models (by robust accuracy) on the datasets considered in scenarios (a) and (b), mentioned in Table 2. The following abbreviations are used: Scen: scenario considered; Best AT/Cl: clean accuracy for the most robust model (by the means of AutoAttack) on the respective dataset, obtained by adversarial training alone; Best P/Cl: clean accuracy for the most robust model (by the means of AutoAttack) on the respective dataset, obtained by adversarial purification alone.
Scen. Dataset Best AT/Cl Best P/Cl (a) Cifar-10 0.9368 0.9082 (b) Cifar-100 0.7522 0.6973

### v6

@@ 6 Conclusion
In this work, we presented a novel adversarial defence mechanism tightly integrating input purification, and classification by an adversarially-trained model – in the form of representation-conditional data purification, followed by a specific logit aggregation. Our method is able to improve upon the current state-of-the-art in Cifar-10, Cifar-100, and TinyImageNet $\ell_{\infty}$ robust classification, w.r.t. both adversarial training and purification approaches alone.
@@ Acknowledgements
Appendix A Justification of Adversarially-balanced batches
@@
Appendix B Justification of the robust aggregation strategy
@@ B.1 Heuristic analysis
In the former scenario, the product of $e^{{l_{i}^{\alpha}}}$ terms can be arbitrarily deflated (up to zero) by lowering the $l_{i^{\star}}^{\alpha^{\star}}$ logit only. In the latter scenario, the logit can be arbitrarily inflated, and such effect is only partially suppressed by normalisation by $Z$ (a sum of $\nicefrac{{1}}{{N}}$-exponentiated terms, $N\geq 1$).
@@
In the former scenario, the effect of logit deflation on a single product term is bounded by $e^{e^{l_{i^{\star}}^{\alpha^{\star}}}}>1$, thus exerting only a minimal collateral effect on the product, through a decrease of $Z$. This effectively prevents aggregation takeover by logit deflation. Similarly to logit averaging, in the latter scenario, the logit can be arbitrarily inflated. However, in this case, the effect of normalisation by $Z$ is much stronger, given its increased magnitude (addends are not $\nicefrac{{1}}{{N}}$-exponentiated, $N\geq 1$).
@@
B.2 Experimental analysis
To further corroborate the choice of the specific aggregation function described in subsection 4.5, we repeat the assessment of Carso under the same conditions described in section 5, the only difference being the use of the alternative aggregation functions analysed in subsection B.1 (i.e. logit and probability averaging aggregation). Results, in terms of both clean and robust accuracy, are shown in Table 4.
@@ B.2 Experimental analysis
In Table 5, we additionally provide the same clean and robust accuracy assessment for the naive and non algorithmically-differentiable majority voting aggregation strategy. In such regard, it is important to remark that the non-differentiability of majority voting results in the vast portion ($99\%$) of gradient samples used by AutoAttack being either zero or not-a-number. Thus, the result of robustness assessment has to be considered unreliable – and not the mark of exceptional robustness – as almost exclusively the effect of gradient obfuscation.
Table 5: Clean (results in italic) and adversarial (results in upright) accuracy resulting from the use of the majority vote aggregation strategy within Carso. The following abbreviations are used: Scen: scenario considered; C/Cl (M.V.): clean accuracy of Carso with majority vote aggregation; C/randAA (M.V.): robust accuracy of Carso with majority vote aggregation, assessed by means of AutoAttack for stochastic defences. Almost the entirety of gradient samples computed by AutoAttack has been deemed unreliable by integrated diagnostics, and the robust accuracy results must be considered untrustworthy.
Scen. Dataset C/Cl (M.V.) C/randAA (M.V.) (a) Cifar-10 0.8691 0.8602 (b) Cifar-100 0.6805 0.6698
As we can see, the use of alternative aggregation strategies leads to minimal variations in the clean accuracy attained, whereas the corresponding robust accuracy sharply decreases – in the case of Cifar-10 even below random chance – as the attacks become increasingly effective (or is unreliable, as it is the case for majority voting). Such results strongly corroborate the use of the normalised doubly-exponential logit product proposed as the aggregation strategy in subsection 4.5 and prove its central role in the overall adversarial robustness and reliability of the method.
Appendix C Architectural details and hyperparameters
In the following section, we provide more precise details about the architectures (subsection C.1) and hyperparameters (subsection C.2) used in the experimental phase of our work.
C.1 Architectures
@@ C.2 Hyperparameters
Fgsm Pgd Pgd+EoT Input clipping $[0.0,1.0]$ $[0.0,1.0]$ $[0.0,1.0]$ # of steps 11 $40$ $200$ Step size $\epsilon_{\infty}$ $0.01$ $0.007$ Loss function CCE CCE CCE # of EoT iterations 11 11 2020 Optimiser SGD SGD
@@
All scenarios Sc. (a) Sc. (b) Sc. (c) Optimiser RAdam++Lookahead RAdam $\beta_{1}$ $0.9$ RAdam $\beta_{2}$ $0.999$ RAdam $\epsilon$ $10^{-8}$ RAdam Weight Decay None Lookahead averaging decay $0.8$ Lookahead steps $6$ Initial LR $5\times 10^{-9}$ Loss function CCE Sampled reconstructions per input $8$ Epochs $200$ $200$ $250$ LR warm-up epochs $25$ $25$ $31$ LR plateau epochs $25$ $25$ $31$ LR annealing epochs $150$ $150$ $188$ Plateau LR $0.064$ $0.064$ $0.0128$ Final LR $4.346\times 10^{-4}$ $4.346\times 10^{-4}$ $1.378\times 10^{-4}$ $\beta$ increase initial epoch $25$ $25$ $32$ $\beta$ increase final epoch $34$ $34$ $43$ Batch size $5120$ $2560$ $1024$ Adversarial batch fraction $0.5$ $0.15$ $0.01$
Appendix D Ablation study on the need for adversarial training
In order to determine whether it is necessary to train on adversarial examples each of the constituent parts of Carso, an ablation study is performed. The architecture of Carso provided in subsection 5.1 is compared in terms of clean and robust accuracy with those ablated as follows. A WideResNet-28-10 model is always used as the classifier.
  • Both the initial instance of the classifier (that used to extract the internal representation) and the final (that used to actually perform classification) are trained on clean examples only.
  • The final classifier is trained on clean examples only, whereas the former is adversarially-trained.
  • The initial classifier is trained on clean examples only, whereas the latter is adversarially-trained.
Results of such comparison are shown in Table 17.
Table 17: Results of the ablation study on the architecture of Carso. The Clean Acc. column shows the test-set accuracy on uncorrupted inputs for the specific ablated model; the randAA Acc. column shows the accuracy of the same model on test-set inputs perturbed by means of the version of AutoAttack suitable for stochastic defences. In the Type of ablation column, any entry different from None (original architecture) indicates the type of training used for the first (before the solidus) and the second (after the solidus) classifier following input-to-output flow, within the ablated Carso architecture.
Dataset Type of ablation Clean Acc. randAA Acc. Cifar-10 clean/clean 0.7314 0.7070 AT/clean 0.6743 << 0.7070 clean/AT 0.8892 0.7975 None 0.8686 0.7613 Cifar-100 clean/clean 0.4395 0.4032 AT/clean 0.4373 << 0.4032 clean/AT 0.6876 0.6716 None 0.6806 0.6665 TinyImageNet-200 clean/AT 0.5677 0.5281 None 0.5632 0.5356
As it is possible to see, only the clean/AT ablation provides clean and adversarial accuracies comparable to that of the original Carso architecture – and indeed, it even determines an improvement on the Cifar-10 and Cifar-100 datasets. On the other hand, adversarial training of the former instance of the classifier is necessary to achieve the best robustness results on TinyImageNet-200.
Keeping in mind that the TinyImageNet-200 dataset is the closest representative considered for larger-scale datasets, and that empirical robust accuracy results only constitute an upper bound of maximally attainable robust accuracy, we support the original Carso architecture as the most effective for the achievement of adversarial robustness on generic image classification datasets – even if at the cost of a slight clean accuracy penalty. Since an adversarially-trained classifier would be used nonetheless as the latter, this does not incur in increased training-time computational intensity.
Appendix E Further results and comparisons
@@
The following section contains additional results and comparisons, in the form of tabular data, that may be of interest to the reader.
In particular, Table 19 compares the clean and robust accuracy of Carso with those of the top-5 adversarial training defences according to the RobustBench Croce et al. (2021) leaderboard [^4] At the time of paper review, Git SHA hash: 78fcc9e48a07a861268f295a777b975f25155964. and the top-5 (if available) purification-based defences according to Lee & Kim (2024), for Cifar-10/Cifar-100.
On shared columns, Table 2 can be considered a subset of Table 19. As discussed in subsection 5.2, Carso tops the comparison in terms of adversarial accuracy, while maintaining a clean accuracy comparable with that of some purification-based models.
Table 18, instead, investigates the same comparison across Carso and its classifier model, for adversarially-pretrained networks different from those described in subsection 5.1 – in particular, worse-performing. As it is possible to see, the increase in end-to-end adversarial accuracy determined by Carso does not depend in absolute terms on the quality of the original classifier model employed.
Table 18: Clean (results in italic) and adversarial (results in upright) accuracy for additional models to those described in subsection 5.2. The following abbreviations are used: AT/Cl: clean accuracy for the adversarially-pretrained model used as the classifier, when considered alone; C/Cl: clean accuracy for the Carso architecture; AT/AA: robust accuracy (by the means of AutoAttack) for the adversarially-pretrained model used as the classifier, when considered alone; C/randAA: robust accuracy for the Carso architecture, when attacked end-to-end by AutoAttack for randomised defences.
Dataset Classification model AT/Cl AT/AA C/Cl C/randAA Cifar-10 Rebuffi et al. (2021) 0.8733 0.6075 0.8152 0.7070 Gowal et al. (2020) 0.8948 0.6280 0.8361 0.7335 Cifar-100 Rebuffi et al. (2021) 0.6241 0.3206 0.5723 0.5580
Table 19: Clean (results in italic) and adversarial (results in upright) accuracy for state-of-the-art adversarial defences, compared to Carso. The Robust Acc. column shows the best (i.e. lowest practically achieved) accuracy on test-set adversarial inputs, as obtained by either the original publication introducing the method, evaluation by AutoAttack as shown on the RobustBench leaderboard, evaluation of bespoke adaptive attacks as shown on the RobustBench leaderboard, or (for purification methods) evaluation by Pgd+EoT from Lee & Kim (2024). The Def. Type column indicates whether the defence is based on adversarial training (AT), purification (P), or both (AT+P). Results for Carso are the same as for Table 2.
Dataset Model Architecture Clean Acc. Robust Acc. Def. type Cifar-10 Bartoldson et al. (2024) WideResNet-94-16 0.9368 0.7371 AT Amini et al. (2024) MeanSparse WideResNet-94-16 0.9568 0.7310 AT Peng et al. (2023) RaWideResNet-70-16 0.9311 0.7107 AT Wang et al. (2023) WideResNet-70-16 0.9325 0.7069 AT Bai et al. (2024b) ResNet-152 ++ WideResNet-70-16 0.9519 0.6971 AT Lin et al. (2024) Diffusion-based 0.9082 0.6641 P Lee & Kim (2024) Diffusion-based 0.9053 0.5688 P Nie et al. (2022) Diffusion-based 0.9043 0.5113 P Hill et al. (2021) Energy-based 0.8412 0.5490 P Yoon et al. (2021) Diffusion-based 0.8612 0.3711 P Ours Carso (WideResNet-28-10) 0.8686 0.7613 AT+P Cifar-100 Wang et al. (2023) WideResNet-70-16 0.7522 0.4266 AT Amini et al. (2024) MeanSparse WideResNet-70-16 0.7513 0.4225 AT Bai et al. (2024b) ResNet-152 ++ WideResNet-70-16 0.8308 0.4180 AT Cui et al. (2023) WideResNet-28-10 0.7385 0.3918 AT Bai et al. (2024a) ResNet-152 ++ WideResNet-70-16 + Mixing Net 0.8521 0.3872 AT Lin et al. (2024) Diffusion-based 0.6973 0.4609 P Ours Carso (WideResNet-28-10) 0.6806 0.6665 AT+P